Angle between pair of Staines x^2+2xy-y^2 is?

Grade:12

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
If we compare this with the standard equation:
ax2+2hxy+by2=0
We get, a =1, h = 1, b = -1
Hence using the formula:
\tan \theta = \frac{2\sqrt{h^2 - ab}}{a + b}
we get:
\tan \theta = \frac{2\sqrt{1^2 +1}}{1 -1}=\infty
Hence,
\theta = 90^{0}
Regards
Sumit